# LIME: Localized Image Editing via Attention Regularization in Diffusion Models

### v1

@@ Abstract
Diffusion models (DMs) have gained prominence due to their ability to generate high-quality, varied images, with recent advancements in text-to-image generation. The research focus is now shifting towards the controllability of DMs. A significant challenge within this domain is localized editing, where specific areas of an image are modified without affecting the rest of the content. This paper introduces LIME for localized image editing in diffusion models that do not require user-specified regions of interest (RoI) or additional text input. Our method employs features from pre-trained methods and a simple clustering technique to obtain precise semantic segmentation maps. Then, by leveraging cross-attention maps, it refines these segments for localized edits. Finally, we propose a novel cross-attention regularization technique that penalizes unrelated cross-attention scores in the RoI during the denoising steps, ensuring localized edits. Our approach, without re-training and fine-tuning, consistently improves the performance of existing methods in various editing benchmarks.
@@
(d) Change the plant color to blue.
Figure 1: LIME: Localized IMage Editing. Our method edits an image based on an edit instruction without needing customized datasets or fine-tuning. The four examples are taken from established papers [52, 29, 53] and compare our edits with the respective state-of-the-art models. The addition of LIME improves all models and allows localized edits that preserve the rest of the image untouched.
footnotetext: The project page can be found at https://enisimsar.github.io/LIME/.
@@ 1 Introduction
In this paper, we address the task of text-guided image editing, explicitly focusing on localized editing, which refers to identifying and modifying any region of interest in an image. This is done regardless of its size and based on textual instructions while preserving the context of the surrounding regions. The difficulty arises from the intertwined nature of image representations within these models, where changes intended for one area can inadvertently affect others [17, 27, 53, 6]. Existing methods often depend on additional user input, such as masking the target area, i.e., Region of Interest (RoI), or providing additional text information, e.g., objects of interest, to pinpoint the editing region [2, 8]. However, these approaches introduce complexity and do not guarantee the precision necessary for seamless editing. Figure 1 highlights localized edits without altering the overall image, a balance that current methods have not yet struck. Advancing localized editing to be more intuitive and effective remains a pivotal direction.
@@
We address the challenge of localized image editing by introducing LIME, that leverages pre-trained InstructPix2Pix [6] without the need for additional supervision, user inputs, or model re-training/fine-tuning. Recent studies [49, 34, 44] have demonstrated that diffusion models are capable of encoding semantics within their intermediate features. LIME utilizes those features to identify segments, then extracts RoI by harnessing attention scores derived from instructions. Other research [7, 1] has shown the significant impact of attention-based guidance on the composition of an image. Accordingly, LIME aims to restrict the scope of edits by regularizing attention scores to enable disentangled and localized edits. By improving these two lines of work, LIME not only offers more effective localized editing as shown in Fig. 1 but also demonstrates a notable advancement by quantitatively outperforming current state-of-the-art methods on four different benchmark datasets.
@@
Our pipeline contains two steps. It first finds semantic segments of the input image. This is achieved based on semantic information encoded in intermediate features. Then, we identify the area to be edited by combining the segments with large cross-attention scores toward the edit instruction. Once we isolate the area to be edited, i.e., RoI, the proposed attention regularization technique is applied to the text tokens to selectively target the RoI to ensure that subsequent editing is accurately focused, avoiding unintended changes to other parts of the image. This two-step approach, first refining targeted areas and then editing within the RoI, ensures that our modifications are accurate and contextually coherent, simplifying the editing process while avoiding unintended alterations to the rest of the image.
The core contributions of this study are:
@@
  • We introduce a localized image editing technique that eliminates the need for fine-tuning or re-training, ensuring efficient and precise localized edits.
@@
  • Our approach leverages the pre-trained model’s intermediate features to segment the image and to identify the regions where modifications will be applied.
  • An attention regularization strategy is proposed, which is employed to achieve disentangled and localized edits within the RoI, ensuring contextually coherent edits.
@@
The experimental evaluation demonstrates that our approach outperforms existing methods in localized editing both qualitatively and quantitatively on four benchmark datasets [52, 20, 5, 6].
@@ 2 Related Work
Text-guided image generation.
Text-to-image synthesis significantly advanced thanks to diffusion models that surpassed prior generative adversarial networks (GANs) [16, 37, 51]. Key developments [19, 43, 10] have resulted in diffusion models that generate highly realistic images from textual inputs [31, 36, 40]. Notably, the introduction of latent diffusion models has significantly increased the computational efficiency of previous methods [38].
Image editing with Diffusion Models.
One direction for image editing is utilizing pre-trained diffusion models by first inverting the input image in the latent space and then applying the desired edit by altering the text prompt [30, 17, 46, 27, 8, 20, 32, 45, 48]. For instance, DirectInversion [20] inverts the input image and then applies Prompt2Prompt [17] to obtain the desired edit, but it may lose details of the input image during inversion. DiffEdit [8], on the other hand, matches the differences in predictions for input and output captions to localize the edit yet struggles with complex instructions. It works in the noise space to edit. Another direction for image editing by using instructions is training diffusion models on triplet data, which contains input image, instruction, and desired image [6, 52, 53, 13]. The latest approach, InstructPix2Pix (IP2P) [6] uses a triplet dataset to train a model for editing images by using instructions. It performs better than previous methods but sometimes generates entangled edits. To tackle this problem, HIVE [53] relies on human feedback on edited images to learn what users generally prefer and uses this information to fine-tune IP2P, aiming to align more closely with human expectations. Alternatively, our method leverages the pre-trained IP2P to localize the edit instruction. Then, instead of manipulating the noise space [8, 2, 29], our method employs attention regularization to achieve localized editing, ensuring the edits are restricted within the RoI. The entire process is done without needing additional data, re-training, or fine-tuning.
Semantics in Diffusion Models.
Intermediate features of diffusion models, as explored in studies like [44, 49, 34, 33], have been shown to encode semantic information. Recent research such as LD-ZNet [34] and ODISE [49] leverages intermediate features of these models for training networks for semantic segmentation. Localizing Prompt Mixing (LPM) [33], on the other hand, utilizes clustering on self-attention outputs for segment identification. Motivated by this success, our method leverages pre-trained intermediate features to achieve semantic segmentation and apply localized edits using edit instructions.
@@ 3 Background
Latent Diffusion Models.
Stable Diffusion (SD) [38] is a Latent Diffusion Model (LDM) designed to operate in a compressed latent space. This space is defined at the bottleneck of a pre-trained variational autoencoder (VAE) to enhance computational efficiency. Gaussian noise is introduced into the latent space, generating samples from a latent distribution $z_{t}$. A U-Net-based denoising architecture [10] is then employed for image reconstruction, conditioned on noise input ($z_{t}$) and text conditioning ($c_{T}$). This reconstruction is iteratively applied over multiple time steps, each involving a sequence of self-attention and cross-attention layers. Self-attention layers transform the current noised image representation, while cross-attention layers integrate text conditioning.
@@
Intuitively, $P$ denotes which areas of the input features will be modified in the attention layer. For cross-attention, this is the area of the image that is affected by one of the conditioning text tokens that define $c_{T}$. Beyond these attention maps, our approach also leverages the output of transformer layers, noted as intermediate features $\phi(z_{t})$, which contain rich semantic content, as highlighted in recent studies [44, 49, 34]. In this work, we modify the cross-attention’s $P$ and leverage the intermediate features $\phi(z_{t})$ to localize edits in pre-trained LDMs.
InstructPix2Pix.
Our method relies on InstructPix2Pix (IP2P) [6], an image-to-image transformation network trained for text-conditioned editing. IP2P builds on top of Stable Diffusion and incorporates a bi-conditional framework, which simultaneously leverages an input image $I$, and an accompanying text-based instruction $T$ to steer the synthesis of the image, with the conditioning features being $c_{I}$ for the image and $c_{T}$ for the text. The image generation workflow is modulated through a classifier-free guidance (CFG) strategy [18] that employs two separate coefficients, $s_{T}$ for text condition and $s_{I}$ for image condition. The noise vectors predicted by the learned network $e_{\theta}$, which corresponds to the individual U-Net step, with different sets of inputs, are linearly combined as represented in Eq. 2 to achieve score estimate $\tilde{e}_{\theta}$. Our method utilizes and modifies the processes for the terms with $c_{I}$ in Eq. 2 to apply localized image editing.
@@ 4 Method
We aim to develop a localized editing method for a pre-trained IP2P without re-training or fine-tuning. The proposed method contains two components: (i) edit localization finds the RoI by incorporating the input image and the edit instruction, and (ii) edit application applies the instruction to RoI in a disentangled and localized manner.
@@ 4.1 Edit Localization
Segmentation:
Our study extends the established understanding that intermediate features of diffusion models encode essential semantic information. In contrast to previous methods that build upon Stable Diffusion [44, 34, 49], our approach works on IP2P and focuses on the features conditioned on the original image ($z_{t}$, $c_{I}$, and $\varnothing$) for segmentation as indicated in Eq. 2. Through experimental observation, we show that these features align well with segmentation objectives for editing purposes. To obtain segmentation maps, we extract features from multiple layers of the U-Net architecture, including both down- and up-blocks, to encompass a variety of resolutions and enhance the semantic understanding of the image. Our preference for intermediate features over attention maps is based on their superior capability to encode richer semantic information, as verified by studies such as [44, 49, 34].
We implement a multi-resolution fusion strategy to refine the feature representations within our proposed model. This involves (i) resizing feature maps from various resolutions to a common resolution by applying bi-linear interpolation, (ii) concatenating and normalizing them along the channel dimension, and (iii) finally, applying a clustering method, such as the K-means algorithm, on fused features. We aim to retain each feature set’s rich, descriptive qualities by following these steps. Moreover, each resolution in the U-Net step keeps different granularity of the regions in terms of semantics and sizes.
Figure 2 demonstrates segmentation maps from different resolutions and our proposed fused features. Each resolution captures different semantic components of the image, e.g., field, racket, hat, dress…. Although Resolution 64 can distinguish objects, e.g., skin and outfit, it does not provide consistent segment areas, e.g., two distinct clusters for lines in the field. On the other hand, lower resolutions, Resolution 16 and 32, can capture coarse segments like lines in the field and the racket. Fusing those features from different resolutions yields more robust feature representations, enhancing the segmentation; see Fig. 2 - Ours. For the extraction of intermediate features, we use time steps between 30 and 50 out of 100 steps, as recommended by LD-ZNet [34].
@@
Instruction: Make her outfit black # of clusters: 8
Figure 2: Segmentation and RoI finding. Resolution Xs demonstrates segmentation maps from different resolutions, while Ours shows the segmentation map from our method. For the cross-attention map, the color yellow indicates high probability, and blue dots mark the $100$ pixels with the highest probability. The last image shows the extracted RoI using blue dots and Ours.
Localization:
Upon identifying the segments within the input image, the proposed method identifies the RoI for the edit using cross-attention maps conditioned on the input image and instruction ($z_{t}$, $c_{I}$, and $c_{T}$) as indicated in Eq. 2. These maps have dimensions of $H_{b}\times W_{b}\times D$, where $H_{b}$ and $W_{b}$ represent the height and width of the features of block $b^{th}$ (up and down blocks), respectively, and $D$ denotes the number of text tokens. Following our segmentation strategy, the cross-attention maps are resized to a common resolution, combined among the spatial dimensions, namely $H$ and $W$, and normalized among the token dimension, $D$. After merging attention maps from different resolutions, the method ignores the <start of text>, stop words, and padding tokens to ignore noisy attention values from unrelated parts of the conditioning text and focuses on the remaining tokens to identify the area that is related to the edit instruction. Then, we get the mean attention score among the tokens to generate a final attention map; see Fig. 2 - Attention. Subsequently, the top $100$ pixels, ablated in Tab. 4, marked by highest probability scores, are identified. Then, all segments that overlap at least one of those pixels are combined to obtain the RoI; see Fig. 2 - Ours, Attention, and RoI.
@@ 4.2 Edit Application
Leveraging the strength of the pre-trained models, we introduce a novel localized editing technique within IP2P. This module manipulates attention scores corresponding to the RoI while ensuring the rest of the image remains the same, thus preventing any unintended alterations outside the RoI. Specifically, this procedure uses the terms with $z_{t}$, $c_{I}$, and $c_{T}$ using the notation of Eq. 2.
Token-based cross attention probabilites
@@
Figure 3: Attention Regularization. Our method selectively regularizes unrelated tokens within the RoI, ensuring precise, context-aware edits without the need for additional model training or extra data. After attention regularization, the probabilities for the related tokens are attending the RoI, as illustrated in the second row.
@@
Figure 4: Qualitative Examples. We test our method on different tasks: (a) editing a large segment, (b) altering texture, (c) editing multiple segments, (d) adding, (e) replacing, and (f) removing objects. Examples are taken from established papers [52, 20, 53]. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.
Attention Regularization:
@@
Previous methods [8, 29, 2] use the noise space instead of attention scores. In contrast, our method introduces targeted attention regularization for selectively reducing the influence of unrelated tokens within the RoI during editing. This approach regularizes attention scores for tokens that are unrelated to the editing task, such as <start of text>, padding, and stop words (denoted as $S$). By adjusting the attention scores ($QK^{T}$) within the RoI, we aim to minimize the impact of these unrelated tokens during the softmax normalization process. As a result, the softmax function is more likely to assign higher attention probabilities within the RoI to tokens that align with the editing instructions. This targeted approach ensures that edits are precisely focused on the desired areas, enhancing the accuracy and effectiveness of the edits while preserving the rest. Given the binary mask for RoI $M$, we modify the result of the dot product $QK^{T}$ of cross-attention layers for unrelevant tokens to a regularization version $R(QK^{T},M)$ as follows:
$\begin{split}R(QK^{T},M)=\begin{cases}QK^{T}_{ijt}-\alpha,&\text{if }M_{ij}=1\text{ and }t\in S\\ QK^{T}_{ijt},&\text{otherwise},\end{cases}\end{split}$ (3)
@@
where $\alpha$ is a large value. Intuitively, we prevent unrelated tokens from attending to the RoI, as shown in Fig. 3. In contrast, related tokens will be more likely to be selected in the RoI, leading to more accurate, localized, and focused edits. This method achieves an optimal balance between targeted editing within the intended areas and preserving the surrounding context, thus enhancing the overall effectiveness of the instruction.
@@
By employing this precise regularization technique within the RoI, our method significantly enhances IP2P. It elevates the degree of disentanglement and improves the localization of edits by tapping into the already-learned features of the model. This targeted approach circumvents the need for re-training or fine-tuning, preserving computational resources and time. It harnesses the inherent strength of the pre-trained IP2P features, deploying them in a focused and effective manner. This precision ensures that edits are contained within the intended areas, underpinning the model’s improved capability to execute complex instructions in a localized and controlled way without the necessity for additional rounds of training or fine-tuning.

### v2

@@ Abstract
Diffusion models (DMs) have gained prominence due to their ability to generate high-quality varied images with recent advancements in text-to-image generation. The research focus is now shifting towards the controllability of DMs. A significant challenge within this domain is localized editing, where specific areas of an image are modified without affecting the rest of the content. This paper introduces LIME for localized image editing in diffusion models. LIME does not require user-specified regions of interest (RoI) or additional text input, but rather employs features from pre-trained methods and a straightforward clustering method to obtain precise editing mask. Then, by leveraging cross-attention maps, it refines these segments for finding regions to obtain localized edits. Finally, we propose a novel cross-attention regularization technique that penalizes unrelated cross-attention scores in the RoI during the denoising steps, ensuring localized edits. Our approach, without re-training, fine-tuning and additional user inputs, consistently improves the performance of existing methods in various editing benchmarks. The project page can be found at https://enisimsar.github.io/LIME/.
@@
(d) Change color of carpet to dark blue.
Figure 1: LIME: Localized IMage Editing. LIME edits an image based on an edit instruction without needing customized datasets, fine-tuning, or explicit information about the object of interest. The addition of LIME improves InstructPix2Pix (IP2P) [7] and its fine-tuned version on MagicBrush (MB) [53], human-annotated, dataset and allows localized edits preserving the rest of the image untouched.
@@ 1 Introduction
In this paper, we address the task of text-guided image editing, explicitly focusing on localized editing, which refers to identifying and modifying any region of interest in an image while preserving the context of the surrounding regions. In particular, we focus in the settings where this needs to be done relying only on textual instructions. The difficulty arises from the intertwined nature of image representations within these models, where changes intended for one area can inadvertently affect others [18, 29, 54, 7]. Existing methods often depend on additional user input, such as masking the target area, i.e., Region of Interest (RoI), or providing additional text information, e.g., objects of interest, to pinpoint the editing region [3, 9, 17]. However, these approaches introduce complexity and do not guarantee the precision necessary for seamless editing. Figure 1 highlights localized edits without altering the overall image, a balance that current methods have not yet struck. Advancing localized editing to be more effective remains a challenging research direction especially when minimizing the need for extensive re-training, fine-tuning, or additional user inputs.
@@
We address the challenge of localized image editing by introducing LIME, that leverages pre-trained InstructPix2Pix (IP2P) [7] without the need for additional supervision, user inputs, or model re-training/fine-tuning. Recent studies [51, 37, 46] have demonstrated that diffusion models are capable of encoding semantics within their intermediate features. LIME utilizes those features to identify segments, then extracts RoI by utilizing attention scores derived from instructions. Other research [8, 1, 30] has shown the significant impact of attention-based guidance on image generation tasks. Accordingly, LIME aims to restrict the scope of edits by regularizing attention scores to enable disentangled and localized edits. By incorporating these two lines of work, LIME not only offers more effective localized editing as shown in Fig. 1 but also demonstrates a notable advancement. This advancement is quantitatively measured by outperforming current state-of-the-art methods on two different benchmark datasets. While we primarily benchmark against IP2P, the method’s principles can be adapted to other relevant approaches, see Supplementary Material.
@@
Our pipeline contains two steps. (i) It first finds semantic segments of the input image. This is achieved based on semantic information encoded in intermediate features. Then, we identify the area to be edited by combining the segments with significant cross-attention scores toward the edit instruction. (ii) Once we isolate the area to be edited, i.e., RoI, the proposed attention regularization is applied to the text tokens cross-attention scores to ensure that the edit is focused to the RoI, avoiding unintended changes to other parts of the image. This two-step approach, first identifying targeted areas and then editing within them, ensures that the edits are local and contextually coherent, simplifying the editing process while avoiding unintended alterations to the rest of the image. The core contributions of this study:
@@
  • We introduce a localized image editing method, LIME, that eliminates the need for fine-tuning, re-training or additional user inputs, ensuring efficient and precise localized edits.
  • We propose a novel attention regularization strategy to ensure that our edits are not only localized but also contextually coherent within the RoI, resulting in seamless and targeted image modifications.
@@
  • Our approach leverages the pre-trained instruction-based editing model’s intermediate features to segment the image and to identify the regions where modifications will be applied.
@@
The experimental evaluation demonstrates LIME outperforms existing methods in localized editing both qualitatively and quantitatively on two benchmark datasets [53, 7].
@@ 2 Related Work
Image editing with Diffusion Models. One direction for image editing is utilizing pre-trained diffusion models by first inverting the input image in the latent space and then applying the desired edit by altering the text prompt [33, 18, 48, 29, 9, 21, 35, 47, 50]. For instance, the input image can be inverted to the diffusion model latent space via Null-text inversion [33], PnPInversion [21] or any other inversion methods and then Prompt-to-Prompt [18] can be applied to obtain the desired edit. Prompt-to-Prompt also introduces an attention modulation technique to control the edit by globally modifying the attention maps of specific tokens. In contrast, the attention regularization of LIME relies on the RoI and thereby is local in the image space. We show that our attention regularization is applicable to Prompt-to-Prompt for improved localized edits in Supplementary Material. DiffEdit [9], on the other hand, matches the differences in predictions for input and output captions that are initialized from the inversion to localize the edit yet it struggles with complex instructions. Moreover, edits are performed by interpolations in the noise space instead of the attention modulation we propose. The key challenge of these methods is the inversion of the real image because it may lose details of the input image.
On the other hand, another direction for image editing by using instructions that eliminates the need for inversion is training diffusion models on triplet data, which contains input image, instruction, and edited image [7, 53, 54, 14]. The most common approach, IP2P [7], performs better than previous methods but sometimes generates entangled edits. To tackle this problem, MagicBrush (IP2P w/MB) [53] fine-tunes IP2P on a more localized human-annotated dataset, and HIVE [54] relies on human feedback on edited images to learn what users generally prefer and uses this information to fine-tune IP2P, aiming to align more closely with human expectations. Concurrently to our work, [17] proposed addressing the same entanglement problem with localized image editing by using grounding properties of IP2P. However, compared to LIME, it requires an additional user input, e.g., the object of interest to edit. Moreover, our method performs edits via attention regularization rather than manipulating the noise space [9, 3, 32].
Semantics in Diffusion Models. Intermediate features of diffusion models, as explored in studies like [46, 51, 37, 36], have been shown to encode semantic information. Recent research such as LD-ZNet [37] and ODISE [51] leverages intermediate features of these models for training networks for semantic segmentation. Localizing Prompt Mixing (LPM) [36], on the other hand, utilizes clustering on self-attention outputs for segment identification. Motivated by this success, our method leverages pre-trained intermediate features to achieve semantic segmentation and apply localized edits using edit instructions.
@@ 3 Background
Latent Diffusion Models. Stable Diffusion (SD) [41] is a Latent Diffusion Model (LDM) designed to operate in a compressed latent space. This space is defined by a pre-trained variational autoencoder (VAE) to enhance computational efficiency. Gaussian noise is introduced into the latent space, generating samples from a latent distribution $z_{t}$. A U-Net-based denoising architecture [11] is then employed for image reconstruction, conditioned on noise input ($z_{t}$) and text conditioning ($c_{T}$). This reconstruction is iteratively applied over multiple time steps, each involving a sequence of self-attention and cross-attention layers. Self-attention layers transform the current noised image representation, while cross-attention layers integrate text conditioning.
@@
Intuitively, $P$ denotes which areas of the input features will be modified in the attention layer. For cross-attention, it corresponds to the affected area by one of the conditioning text tokens that define $c_{T}$. Beyond the attention maps, our approach leverages the output of transformer layers, noted as intermediate features $\phi(z_{t})$, which contain rich semantic content, as highlighted in recent studies [46, 51, 37]. In this work, we modify the cross-attention’s $P$ and leverage the intermediate features $\phi(z_{t})$ to localize edits.
InstructPix2Pix. Our method relies on IP2P [7], an image-to-image transformation network trained for text-conditioned editing. IP2P builds on top of Stable Diffusion and incorporates a bi-conditional framework, which simultaneously leverages an input image $I$, and an accompanying text-based instruction $T$ to steer the synthesis of the image, with the conditioning features being $c_{I}$ for the image and $c_{T}$ for the text. The image generation workflow is modulated through a classifier-free guidance (CFG) strategy [20] that employs two separate coefficients, $s_{T}$ for text condition and $s_{I}$ for image condition. The noise vectors predicted by the learned network $e_{\theta}$, which corresponds to the individual U-Net step, with different sets of inputs, are linearly combined as represented in Eq. 2 to achieve score estimate $\tilde{e}_{\theta}$. Our method utilizes and modifies the processes for the terms with $c_{I}$ in Eq. 2 to apply localized image editing.
@@ 4 Method
We aim to develop a localized editing method for a pre-trained IP2P without re-training, fine-tuning or the need for additional user inputs besides an edit prompt. The proposed method contains two components: (i) Edit Localization finds the RoI by using the input image and the edit instruction, see Sec. 4.1, and (ii) Edit Application applies the instruction to RoI in a disentangled and localized manner, see Sec. 4.2.
@@ 4.1 Edit Localization
Segmentation Our study extends the established understanding that intermediate features of diffusion models encode essential semantic information. In contrast to previous methods that build upon Stable Diffusion [46, 37, 51], our approach works on IP2P and focuses on the features conditioned on the original image ($z_{t}$, $c_{I}$, and $\varnothing$) for segmentation as indicated in Eq. 2. Through experimental observation, we show that these features align well with segmentation objectives for editing purposes. To obtain segmentation maps, we extract features from multiple layers of the U-Net architecture, including both down- and up-blocks, to encompass a variety of resolutions and enhance the semantic understanding.
We implement a multi-resolution fusion strategy to refine feature representations within our proposed model and cluster them to extract segments. This involves (i) resizing feature maps from various resolutions to a common resolution by applying bi-linear interpolation, (ii) concatenating and normalizing them along the channel dimension, and (iii) finally, applying a clustering method, such as the K-means algorithm, on fused features. Following these steps, we aim to retain each feature set’s rich, descriptive qualities. Moreover, each resolution in the U-Net step maintains a different granularity of the regions regarding semantics and sizes.
Figure 2 demonstrates the resulting segmentation maps from different resolutions and our proposed fused features. Each resolution captures different semantic components of the image (e.g., field, racket, hat and dress). Although Resolution 64 can distinguish objects (e.g., skin and outfit), it does not provide consistent segment areas, e.g., two distinct clusters for lines in the field. On the other hand, lower resolutions, Resolution 16 and 32, capture coarse segments like lines in the field and the racket. Fusing those features from different resolutions yields more robust features, enhancing the segmentation; see Fig. 2 - Ours. For the extraction of intermediate features, we use time steps between 30 and 50 out of 100 steps, as recommended by LD-ZNet [37].
@@
Instruction: Make her outfit black
# of clusters: 8
Figure 2: Segmentation and RoI finding. Resolution Xs demonstrates segmentation maps from different resolutions, while Ours shows the segmentation map from our multi-resolution fusion method explained above. For the cross-attention map, the color yellow indicates high probability, and blue dots mark the N points with the highest probability. The last image shows the extracted RoI using blue dots and Ours.
Localization Upon identifying the segments within the input image, the proposed method identifies the RoI for the edit using cross-attention maps conditioned on the input image and instruction ($z_{t}$, $c_{I}$, and $c_{T}$) as indicated in Eq. 2. These maps have dimensions of $H_{b}\times W_{b}\times D$, where $H_{b}$ and $W_{b}$ represent the height and width of the features of block $b^{th}$ (up and down blocks), respectively, and $D$ denotes the number of text tokens. Following the same strategy for segmentation, the cross-attention maps are resized to a common resolution, which is $64\times 64$, combined among the spatial dimensions, namely $H$ and $W$, and normalized among the token dimension, $D$. Then, our method discards unrelated tokens, e.g., padding, stop words defined by NLTK [6], and <start of text> (SoT) defined by CLIP [39]. SoT is an unrelated token, as recommended by [8], while <end of text> (EoT) is a related token since it is used as the feature representation of the text [39]. We verify experimentally that filtering these tokens allows LIME to focus on related tokens, accurately identifying the RoI in accordance with the edit instruction. Then, we get the mean attention score among the tokens to generate a final attention map; see Fig. 2 - Attention, that should correlate with the area of the image where the edit will be applied. Subsequently, we identify the top N points with the highest attention score, e.g., $N=100$ points - approximately $2.5\%$ of $64\times 64$ attention maps, ablated in Tab. 2. Eventually, all segments that overlap at least one of those points are combined to obtain the RoI; see Fig. 2 - Ours, Attention, and RoI.
@@ 4.2 Edit Application
This component proposes a novel attention regularization which manipulates attention scores corresponding to the RoI while ensuring the rest remains the same, thus preventing any unintended alterations outside the RoI, i.e., localized image editing. Specifically, this procedure uses the terms with $z_{t}$, $c_{I}$, and $c_{T}$ using the notation of Eq. 2.
Token-based cross-attention probabilities
@@
Figure 3: Attention Regularization. Our method selectively regularizes unrelated tokens (SoT and stop words: her) within the RoI, ensuring precise, context-aware edits without the need for extra model training or user inputs. After attention regularization, the probabilities for the related tokens are attending the RoI, as illustrated in the second row.
@@
Figure 4: Qualitative Examples. We test our method on different tasks: (a) editing a large segment, (b) altering texture, (c) editing multiple segments, (d) adding, (e) replacing, and (f) removing objects. The integration of LIME enhances the performance of baselines, enabling localized edits while preserving the remaining image areas.
@@
Attention Regularization Previous methods [9, 32, 3] use the noise space instead of attention scores. In contrast, our method introduces targeted attention regularization for selectively reducing the influence of unrelated tokens within the RoI during the editing process. This approach regularizes attention scores for tokens that are unrelated to the editing task, such as SoT, padding, and stop words, defined in Sec. 4.1 (denoted as $S$). By adjusting the attention scores ($QK^{T}$) within the RoI, we aim to minimize the impact of these unrelated tokens during the softmax normalization process. As a result, the softmax function is more likely to assign higher attention probabilities within the RoI to tokens that align with the editing instructions, i.e., related tokens. This targeted approach ensures that edits are precisely focused on the desired areas, enhancing the accuracy and effectiveness of the edits while preserving the rest. Given the binary mask for RoI $M$, we modify the result of the dot product $QK^{T}$ of cross-attention layers for unrelated tokens to a regularization version $R(QK^{T},M)$ as follows:
$\begin{split}R(QK^{T},M)=\begin{cases}QK^{T}_{ijt}-\infty,&\text{if }M_{ij}=1\text{ and }t\in S\\ QK^{T}_{ijt},&\text{otherwise}.\end{cases}\end{split}$ (3)
@@
Intuitively, we prevent unrelated tokens from attending to the RoI, while related tokens will be more likely to be selected in the RoI, leading to more accurate, localized, and focused edits, as shown in Fig. 3. This method achieves an optimal balance between targeted editing within the intended areas and preserving the surrounding context, thus enhancing the overall effectiveness of the instruction.
@@
By employing our regularization technique within the RoI significantly enhances IP2P. It elevates the degree of disentanglement and improves the localization of edits by tapping into the already-learned features of the model. This targeted approach circumvents the need for re-training or fine-tuning , preserving computational resources and time. It harnesses the inherent strength of the pre-trained IP2P features, deploying them in a focused and effective manner. This precision ensures that edits are contained within the intended areas, underpinning the model’s improved capability to execute complex instructions in a localized and controlled way without the necessity for additional rounds of training, fine-tuning or additional user inputs.
5 Experiments
To assess the effectiveness of our model, we integrate it into the widely recognized IP2P [7], referenced as IP2P + LIME throughout this section. The original IP2P model serves as our standard benchmark for all analyses. Our evaluation involves a comprehensive quantitative comparison of our approach against other state-of-the-art methods, utilizing established datasets and metrics, complemented by a user study. While LIME can be integrated with other models [18, 54], we limit the experimental analysis in the main paper to only IP2P for clarity. Additional results with other models will be provided in Supplementary Material.
Input Image
IP2P
+ LIME
IP2P w/MB
+ LIME
Change the alcohol shelf into a bookshelf.
Change the color of the lighthouse into completely red.
Put Stone Henge as the background of the scene.
Remove the forks from the shelf.
Change the robin to a silver robin sculpture.
Figure 5: More Qualitative Examples. We test our method on different tasks: (a) modifying multiple objects, (b) changing color, (c) adding, (d) removing objects, and (e) changing texture. The integration of LIME enhances the performance of all models, enabling localized edits while maintaining the integrity of the remaining image areas.
5.1 Evaluation Datasets and Metrics
MagicBrush [53] Its test split offers an evaluation pipeline with 535 sessions - refer to the source images used for iterative editing instructions - and 1053 turns - denote the individual editing steps within each session. It employs L1 and L2 norms to measure pixel accuracy, CLIP-I, and DINO embeddings for assessing image quality via cosine similarity, and CLIP-T to ensure that the generated images align accurately with local textual descriptions.